A company is using a product with a daily average demand of 500 units and a daily standard deviation of 50. The ordering lead time for this product is 9 business days and the targeted service level is 95%. The reorder point for this product is approximately Select one:

a. 3785 units.

b. 4500 units.

c. 4627units.

d. 4747units.
Answer #1

Z value of 95% of service level = NORMSINV ( 0.95) = 1.6448

The formula for reorder point

= Daily demand x Lead time ( days) + Z value x standard deviation x square root ( lead time )

= 500 x 9 + 1.6448 x 50 x square root ( 9)

= 4500 + 1.6448 x 50 x 3

= 4500 + 246.72

= 4746.72 ( 4747 rounded to nearest whole number )

ANSWER : 4747